NvidiaIntely (nvidiaintely):

Given 4.8 moles of a gas at 37 degrees Celsius and at 792 torr, what is the volume of the gas? (The ideal gas constant is 0.0821 L · atm/mol · K and 1 atm = 760 torr.) 1.40 x 101 L 1.84 x 10-2 L 1.17 x 102 L 1.54 x 10-1 L

OpenStudy (helpz-q):

So first tell me what you think it is... explain your thoughts as well as opinions

OpenStudy (helpz-q):

And actually i am a 9th grade teacher and i have been teaching my class this for a while now

OpenStudy (jiteshmeghwal9):

p=pressure=792/760 atm n=4.8 T=37+273 'kelvin R=0.0821

OpenStudy (jiteshmeghwal9):

v=nRT/P

NvidiaIntely (nvidiaintely):

I think its option C.

OpenStudy (francesca_marie):

It's all in the equation: PV = nRT You know the following: P = 792 torr T = 310 degrees K (you must convert to degrees K) n = 4.8 moles So rearrange to solve for V V = nRT/V Now, convert the P into atm (divide 790 by 760) and plug in the value of R to get your volume, in L. or choose a value for R that is in units to L-torr/deg mol (this is 62.364) and solve it that way. In either case, I get 118, which appears to be answer C

OpenStudy (jiteshmeghwal9):

yes u r correct @NvidiaIntely

NvidiaIntely (nvidiaintely):

Thanks. BTW @Francesca_Marie, you got that yahoo. I was just reading that....